# Task-Oriented Automatic Fact-Checking with Frame-Semantics

### v2

@@ Abstract
We propose a novel paradigm for automatic fact-checking that leverages frame semantics to enhance the structured understanding of claims and guide the process of fact-checking them. To support this, we introduce a pilot dataset of real-world claims extracted from PolitiFact, specifically annotated for large-scale structured data. This dataset underpins two case studies: the first investigates voting-related claims using the Vote semantic frame, while the second explores various semantic frames based on data sources from the Organisation for Economic Co-operation and Development (OECD). Our findings demonstrate the effectiveness of frame semantics in improving evidence retrieval and explainability for fact-checking. Finally, we conducted a survey of frames evoked in fact-checked claims, identifying high-impact frames to guide future work in this direction.
@@ 1 Introduction
However, these studies only consider data which have already been processed and prepared for easy consumption by readers. This reliance on highly-curated data hinders their ability to fact-check novel claims which do not have such clean evidence readily available. To bridge this gap our work proposes a novel pilot dataset of real-world factual claims related to two high-volume structured databases on U.S. congressional voting records and country statistics.color=yellow]Saying we created the dataset to fill the gap gives me the feeling of we invent the nail. Instead, consider saying the verification of many claims rely on data that only exist in much larger tables. These databases significantly differ from past works in both volume and detail of their tables. For example, Wikipedia tables consist of an average of 13 rows Chen et al. (2020) while our database of country statistics has an average of 596,000 rows across over 400 tables. This scale of data is incompatible with most previous methods, which often rely on large language models (LLMs), as the tables themselves are too large to be analyzed, even for those solutions designed for large tables Nahid and Rafiei (2024).
@@
Fact-checking novel claims requires fine-grained methodology to understand the claim and to map important parts of the claim to their relevant data. Furthermore, transparency and explainability are critical in fact-checking. We identify frame semantics Baker et al. (1998) as a viable solution which addresses each of these points. Frame semantics is a linguistic framework that explores how meaning is encoded through structured representations called frames. These frames capture the important elements and relationships in various situations and enable explainable understanding of claims and extraction of their key parts. The existence of specific frames in a claim also enables fine-grained task-specific handling of claims for fact-checking.
@@
Our proposed pilot dataset serves as the foundation for two case studies that demonstrate how frame semantics can be used to guide specific fact-checking processes and enhance the explainability of automatic fact-checking. In the first case study, we focus on the Vote frame Arslan et al. (2020), which models how an Agent (e.g., a legislator) interacts with an Issue (e.g., a bill) through the act of voting and optionally takes a Position (for or against) on the Issue. These Issue and Agent elements are then used to implement voting-specific fact-checking methods. The second case study follows similar principles on a much more diverse set of frames applied to a high-volume, structured database produced from hundreds of datasets collected from the Organisation for Economic Co-operation and Development (OECD).
To evaluate these case studies, we collected 79 and 68 real-world claims for the voting records and OECD statistics, respectively. Each claim is manually annotated for its particular frames and fact-check verdict retrieved from PolitiFact.
@@
Our case studies demonstrate that frame semantics offers a valuable framework for task-specific fact-checking. Frame-semantics can also offer additional performance benefits for particular tasks. For example, on Vote and OECD-related claims, retrieving relevant evidence by querying with extracted task-specific frame elements outperformed using the entire claim by 14% and 11%, respectively. The performance of particular retrieval methodologies also significantly varied between the two case studies, indicating a need to select different approaches for each specific task to achieve optimal performance. Furthermore, instructing large language models (LLMs) to fact-check claims benefits from different prompts depending on the type of claim. These factors further showcase the benefit of a task-specific approach to automatic fact-checking compared to a suboptimal generalized approach.
@@
Finally, to understand which frames are most commonly evoked in factual claims, we surveyed claims fact-checked by PolitiFact and identified a heavy skew toward a few highly frequent frames. Notably, the frames used in our case studies are among the most frequently evoked frames in PolitiFact fact-checks. This alignment highlights the practical relevance of our approach, as it directly addresses the semantic structures most commonly encountered in professional fact-checking. color=yellow]This paragraph is almost like after-thought. need to present it more naturally and perhaps in better places. — doesn’t really feel like there is a better location.
@@
  • We developed a pilot dataset for automatic fact-checking of real-world claims using high-volume structured data. [^1] https://github.com/anon-sub-repo/fact-checker
  • We conducted a novel survey of frames evoked in PolitiFact fact-checks (Table 2), enabling targeting high-impact frames for future studies.
  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo and source code. [^2] https://anon-sub-repo.github.io/
@@ 2 Related Work
Previous studies Aly et al. (2021); Gupta et al. (2020); Chen et al. (2020) often use the term “large-scale datasets” to refer to collections with a large quantity of claims and tables. In these cases, the emphasis is on the breadth of the dataset, with a significant number of claims and corresponding tables, typically featuring a modest number of rows per table. For instance, FEVEROUS Aly et al. (2021) focuses on fact-checking against Wikipedia tables, which average around 13 rows per table, enabling techniques that rely on lightweight parsing and exhaustive search methods to handle the smaller, well-curated tabular structures. In contrast, our work addresses high-volume data, where individual tables can contain hundreds of thousands or even millions of rows, as seen in the OECD database used in our case studies, which averages 596,000 rows per table. This shift in scale introduces unique challenges for fact-checking systems, particularly in evidence retrieval.
@@
For structured data, past work has studied using fine-tuning methods Zhao et al. (2022); Gu et al. (2022); Jiang et al. (2022) and LLM-based methods Ye et al. (2023); Chen et al. (2021); Cheng et al. (2023) for fact verification. Dater Ye et al. (2023) is a fact-verification system which simplifies the fact-checking process by decomposing claims into sub-questions using LLMs. It is the best-performing model on the TabFact dataset Chen et al. (2020). Our work does not focus on fact verification; however, we choose to utilize LLM-based methods for our fact-verification step as they do not require retraining the model when new data is used.
@@
Madabushi (2024) compares search methods for finding textual entailment facts in RAG-based systems, including a frame-based method. The authors found that their frame-based approach significantly improved recall@K compared to other search-based methods. Additionally, many previous studies have used frame semantics, or similar semantic parsing methods, in creating task-oriented systems in other domains Neves Ribeiro et al. (2023); Gupta et al. (2018); Chen et al. (2014).
Figure 1: An example of our proposed paradigm. First, the frame and frame elements (FEs) are extracted using a frame-semantic parser (FSP). Then, FEs are mapped to a table which can then be queried to gather evidence for the claim. Finally, the evidence and claim are passed into a fact verification model to check the truthfulness of the claim.
@@ 3 Fact-Checking Paradigm
We break down the task of fact-checking into three key steps: claim understanding, evidence retrieval, and claim-evidence alignment. Figure 1 shows an example of how our system processes the statement “Japan has the 2nd highest life expectancy.” color=yellow]Make it clear in the figure that “Occupy Rank” is the frame. Also mention it in the figure caption. color=yellow]Why is “life-expectancy” schema listed twice, with slight difference (w/ and w/o measure)? color=yellow]In figure 1, spell out FSP as “Frame Semantic Parsing (FSP)” color=yellow]In Figure 1, the yellow boxes are procedures. Other colors are data/concept. But they are all represented as boxes. In addition to color, can we further differentiate them using something else? color=yellow]Figure 1 and this section may appear inconsistent. “claim understanding”, “evidence retrieval”, “claim-evidence aligenment” is no where to be found in the figure. On the other hand, isn’t not clear table mapping, table querying, FSP, fact verfication belong to which of the 3 steps. It is like we have presented two alternative naming systems or workflows.
@@ 3.1 Claim Understanding
In the first step, the system focuses on understanding the specifics of the claim. Frame-semantic parsing Das et al. (2014) enables the extraction of frames and their frame elements from claims.color=green]can make the concept of FSP more clear. Even the next sentence (“identifying frames and … FEs …”) kind of defines it and can be used here instead. color=green]Need references for FSP By extracting this representation, the system gains a comprehensive understanding of the claim, which is critical for downstream tasks like evidence retrieval and verification. For example, in Figure 1, the Occupy_Rank frame is evoked, helping the system understand the claim in terms of the Rank of an Item along a Dimension.color=green]Be consistent. The frame name in the figure doesn’t use underscore.
Frame-semantics not only facilitates claim understanding but also guide the fact-checking process by linking frame elements to appropriate data sources. While some frame elements, such as the Agent and Issue in the Vote frame, map directly to specific database tables (i.e., congress members and bills), other frames may require a predicted mapping based on the text. Figure 1 illustrates an example of this predicted mapping.color=yellow]we can find ways to articulate direct vs. predicted mapping more clearly. Probably won’t have time before deadline.
@@ 3.2 Data Collection and Evidence Retrieval
Evidence retrieval is guided by the frame elements identified in the claim understanding phase. These frame elements act as filtering conditions for querying the appropriate tables or data sources. For example, in Figure 1, the Dimension frame element is used to query the relevant table for life expectancy statistics. By focusing on specific spans of the claim, this method minimizes confusion and enhances retrieval accuracy. Additionally, due to the fine-grained control enabled by frame-semantics, our proposed system can support unstructured data by querying large text corpora instead of a database. color=yellow]If we don’t have experiments to support this, we don’t want to discuss it here. A future work or discussion section at the end is the right place.
3.3 Claim-Evidence Alignment
The final step involves aligning the retrieved evidence with the original claim. Frame elements play a critical role in ensuring that the evidence accurately reflects the claim’s context. By only using specific spans of the claim, the system can align the claim and evidence more precisely. This approach also enhances explainability, as the system can point to the text used for evidence retrieval.
Determining the truthfulness of a claim requires synthesizing retrieved evidence and understanding their connection to the claim as well as the intent behind the claim. For example, in Figure 1, the claim states that Japan has the 2nd best life expectancy, yet the data indicates that it is in fact the best. While this claim is not entirely accurate, its underlying premise holds true—Japan has an exceptionally high life expectancy. As a result, the model would predict it as mostly true. color=yellow]This is an intriguing example. But does our approach tackles such a challenge, do we explain how it does that in principle, and do we have empirical results to demonstrate it? — this consideration is included as part of our prompt, but not explicitly discussed in depth as it would be quite difficult to study and probably be its own paper.
While fine-tuned approaches have shown slightly better performance on table-based fact-checking Ye et al. (2023), we employ an LLM-based approach due to their ability to handle novel inputs without needing to retrain and for their ability to provide natural language explanations of predictions. Because this step is not a key focus of our work, we leave fact verification model-agnostic to allow future improvements. Thus, we can replace the current LLM with any fact verification model which takes structured evidence and a claim to perform fact verification.
@@ 4 Fact-Checking Case Studies
color=yellow]This is really an important ToDo item. Untify the two case studies in the same terminology framework.
color=yellow]Another important ToDo item to strengthen the paper: incorporate more case studies.
@@ 4.1 Voting Records
We compiled a dataset of official U.S. congressional voting records from the Congress Github Repository. [^3] https://github.com/unitedstates/congress The dataset includes 342,466 bills from the 93rd Congress to the 117th Congress, with 7,195,798 individual votes across 22,447 roll calls. Each member of Congress cast an average of 4,230 votes from the 101st to the 117th Congress, and the dataset includes a total of 12,677 unique members of Congress since the 1st Congress.color=yellow]We are mentioning 93rd, 101st, and 1st. If we only have votes since 101st Congress, why did we include 93rd Congress or even the 1st Congress? At least mentioning statistics for them becomes less relevant. — this is the data we have, it wouldn’t really make sense to say we only have data from the 101st since we technically have members from before
@@
To verify voting records, mapping the Agent FE to the correct member of Congress is essential. We use SQL queries to match congress members whose names are similar to the words in the Agent FE (details in Appendix C.4). In cases of name ambiguity we default to the more recent member.
This stage presents several challenges. Claims often use nicknames, such as “Sleepy Joe” for Joe Biden or “Meatball Ron” for Ron DeSantis. To address this, we extracted two lists of political nicknames from Wikipedia Wikipedia contributors (2024a, b) to map nicknames to their corresponding congress members. Similarly, some members go by shortened or preferred names, such as “Joe” instead of Joseph or “Ted” for Rafael Edward Cruz. We handle this by utilizing a list 555 https://github.com/carltonnorthern/nicknames of congress members’ preferred names, supplemented by common alternatives when necessary.
@@
Identifying the correct bill based on the extracted Issue FE is challenging because the Issue FE can refer to various types of information, such as abstract topics (e.g., “gun control”), specific actions or bills (e.g., the “Inflation Reduction Act of 2022”), or outcomes of legislation (e.g., “preventing women from getting abortions”). Additionally, bills often do not include the colloquial terms commonly used to describe them. For example, the “STOP School Violence Act of 2018” may be informally described as expanding access to guns in schools, even though the bill itself does not use this phrasing. Because of these challenges, keyword-based search is insufficient for accurate evidence retrieval. To address this, we employ an asymmetric semantic similarity model, as described in Section 5.4, enabling us to identify bills that are semantically similar to the Issue FE, even when the language used in the claim and the bill differs significantly.
@@
Determining whether a claim is refuted or supported by evidence presents several challenges. First, the system cannot rely solely on the vote (Yea or Nay) and the Position Frame Element (for or against).  [^6] Note that Position FE is extracted from claims by our frame semantic parsing model. However, only Agent and Issue FEs are used for evidence retrieval in voting-related claims. Claims may be made without a Position FE, and the relationship between the vote on a bill and the claim’s Position FE can differ. For instance, a Yea vote on a bill may not indicate support for the claimed Issue. Consider the claim, “DeSantis voted against allowing abortions,” in conjunction with a Yea vote on a bill that bans abortion. Here, the vote supports the claim despite the discrepancy between the Position FE (against) and the vote on the bill (for/yea). Second, assessing whether a claim is supported or refuted by a bill vote requires an understanding of the bill itself and its implications. color=yellow]This analysis is valid, but do we discuss solution for “understanding of the bill itself and its implications”. Reading this makes me feel it is doing hands-waving.— LLM is prompted to consider potential impacts of voting a certain way on a given bill

### v3

@@ Abstract
Automatic fact-checking is a critical tool for addressing the growing challenge of misinformation, particularly when verifying novel claims that lack previously curated evidence. Prior work has largely focused on unstructured or highly-curated data sources, limiting scalability and generalization. In this work, we introduce a novel paradigm for task-oriented automatic fact-checking using frame semantics to improve interpretability and evidence retrieval from high-volume structured datasets. We present a new pilot dataset of real-world factual claims grounded in two large-scale databases: U.S. congressional voting records and OECD country statistics. Across two case studies using these datasets, we demonstrate how frame elements can guide fine-grained retrieval and automatic fact-checking. Our experiments show that frame element-driven evidence retrieval improves recall by 14% and 11% over full-claim baselines in the voting and OECD case studies, respectively. We further analyze frame distributions in PolitiFact fact-checked claims and find a strong alignment with the frames targeted in our study. Overall, our results highlight frame semantics as a promising foundation for scalable, interpretable, and domain-adaptable automatic fact-checking.
@@ 1 Introduction
However, these studies only consider data which have already been processed and prepared for easy consumption by readers. This reliance on highly-curated data hinders their ability to fact-check novel claims which do not have such clean evidence readily available. To bridge this gap our work proposes a novel pilot dataset of real-world factual claims related to two high-volume structured databases on U.S. congressional voting records and country statistics. These databases significantly differ from past works in both volume and detail of their tables. For example, Wikipedia tables consist of an average of 13 rows Chen et al. (2020) while our database of country statistics has an average of 596,000 rows across over 400 tables. This scale of data is incompatible with most previous methods, which often rely on large language models (LLMs), as the tables themselves are too large to be analyzed due to context length limitations, even for those solutions designed for large tables Nahid and Rafiei (2024).
@@
Fact-checking novel claims requires fine-grained methodology to understand the claim and to map important parts of the claim to their relevant data. Furthermore, transparency and explainability are critical in fact-checking. We identify frame semantics Baker et al. (1998) as a viable solution which addresses each of these points. Frame semantics is a linguistic framework that explores how meaning is encoded through structured representations called frames. These frames capture the essential elements and relationships of the situations described by claims, supporting explainable understanding and extraction of key components from the claims. The existence of specific frames in a claim also enables fine-grained task-specific handling of the claim for fact-checking. In our paradigm, we use the frames evoked, and manually-selected frame elements, in a given claim to identify relevant database tables for automatically fact-checking it.
@@
Our proposed pilot dataset serves as the foundation for two case studies that demonstrate how frame semantics can be used to guide specific fact-checking processes and enhance the explainability of automatic fact-checking. In the first case study, we focus on the Vote frame Arslan et al. (2020), which models how an Agent (e.g., a legislator) interacts with an Issue (e.g., a bill) through the act of voting and optionally takes a Position (for or against) on the Issue. These Issue and Agent elements are then used to implement voting-specific fact-checking methods. The second case study follows similar principles on a much more diverse set of frames applied to a high-volume, highly diverse structured database produced from hundreds of datasets collected from the Organisation for Economic Co-operation and Development (OECD).
@@
To evaluate these case studies, we collected 79 and 68 real-world claims for the voting records and OECD statistics, respectively. Each claim is manually annotated for its particular frames and fact-check verdict retrieved from PolitiFact. Our case studies demonstrate that frame semantics offers a valuable framework for task-specific fact-checking. Frame-semantics can also offer additional performance benefits for particular tasks. For example, on Vote and OECD-related claims, retrieving relevant evidence by querying with extracted task-specific frame elements outperformed using the entire claim by 14% and 11%, respectively. The performance of particular retrieval methodologies also significantly varied between the two case studies, indicating a need to select different approaches for each specific task to achieve optimal performance. Furthermore, instructing large language models (LLMs) to fact-check claims benefits from different prompts depending on the type of claim. These factors further showcase the benefit of a task-specific approach to automatic fact-checking compared to a suboptimal generalized approach.
@@
Finally, to understand which frames are most commonly evoked in factual claims, we surveyed claims fact-checked by PolitiFact and identified a heavy skew toward a few highly frequent frames. Notably, the frames used in our case studies are among the most frequently evoked frames in PolitiFact fact-checks. This alignment highlights the practical relevance of our approach, as it directly addresses the semantic structures most commonly encountered in professional fact-checking.
@@
  • We developed a pilot dataset for automatic fact-checking of real-world claims using high-volume structured data and released it along with our source code. [^1] https://github.com/idirlab/claimlens-case-studies
  • We conducted a novel survey of frames evoked in PolitiFact fact-checks, enabling targeting high-impact frames for future studies.
  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo. [^2] https://idir.uta.edu/claimlens/
@@ 2 Related Work
Previous studies Aly et al. (2021); Gupta et al. (2020); Chen et al. (2020) often use the term “large-scale datasets” to refer to collections with a large quantity of claims and tables. In these cases, the emphasis is on the breadth of the dataset, with a significant number of claims and corresponding tables, typically featuring a modest number of rows per table. For instance, FEVEROUS Aly et al. (2021) focuses on fact-checking against Wikipedia tables, which average around 13 rows per table, enabling techniques that rely on lightweight parsing and exhaustive search methods to handle the smaller, well-curated tables.
However, a key limitation of these datasets is their reliance on curated data, largely extracted from Wikipedia. Wikipedia tables, while useful for research, often represent a simplified or filtered version of underlying data from primary sources. This includes implicit decisions about which metrics to include (e.g., GDP per capita in USD vs. PPP), how to aggregate regional data, or how to normalize column formats for readability. In contrast, our work avoids these curated intermediaries and instead operates directly on primary source data in its original, often more complex format. For example, the OECD tables used in our study contain nearly 600,000 rows per table on average.
This shift from curated to raw primary data introduces new challenges, particularly in evidence retrieval and reasoning, as systems must handle large-scale, high-dimensional tables without relying on pre-filtered or human-edited summaries. It also better reflects real-world fact-checking settings, where claims must be verified against the same raw data sources used by policymakers or researchers.
@@
For structured data, past work has studied using fine-tuning methods Zhao et al. (2022); Gu et al. (2022); Jiang et al. (2022) and LLM-based methods Ye et al. (2023); Chen et al. (2021); Cheng et al. (2023) for fact verification. Dater Ye et al. (2023) is a fact-verification system which simplifies the fact-checking process by decomposing claims into sub-questions using LLMs. It is the best-performing model on the TabFact dataset Chen et al. (2020). Our work does not primarily focus on fact verification; rather, our main emphasis is on evidence retrieval. Nevertheless, we employ LLM-based methods for the fact-verification step, as they offer the practical advantage of not requiring model retraining when new data becomes available.
@@
Madabushi (2024) compares search methods for finding textual entailment facts in RAG-based systems, including a frame-based method. The authors found that their frame-based approach significantly improved recall@K compared to other search-based methods. Additionally, many previous studies have used frame semantics, or similar semantic parsing methods, in creating task-oriented systems in other NLP domains Neves Ribeiro et al. (2023); Gupta et al. (2018); Chen et al. (2014).
Figure 1: An example of our proposed paradigm. First, the frame Occupy Rank and frame elements (FEs) are extracted using a frame-semantic parser (FSP). Then, FEs are mapped to a table which can then be queried to gather evidence for the claim. Finally, the evidence and claim are passed into a fact verification model to check the truthfulness of the claim.
@@ 3 Fact-Checking Paradigm
We break down the task of fact-checking into three key steps: claim understanding, evidence retrieval, and fact verification. Figure 1 shows an example of how our system processes the statement “Japan has the 2nd highest life expectancy.”
@@ 3.1 Claim Understanding
In the first step, the system focuses on understanding the specifics of the claim. Frame-semantic parsing Das et al. (2014) enables the extraction of frames and their frame elements from claims. By extracting this representation, the system gains a comprehensive understanding of the claim, which is critical for downstream tasks such as evidence retrieval and verification. For example, in Figure 1, the Occupy Rank frame is evoked, which characterizes the claim in terms of an Item occupying a certain Rank along a Dimension. This semantic structure provides useful context that can support the system’s interpretation and downstream processing of the claim.
Frame-semantics not only facilitates claim understanding but also guides the fact-checking process by linking frame elements to appropriate data sources. While some frame elements, such as the Agent and Issue in the Vote frame, map directly to specific database tables (i.e., congress members and bills), other frames may require a predicted mapping based on the text. Figure 1 illustrates an example of this predicted mapping.
@@ 3.2 Data Collection and Evidence Retrieval
Evidence retrieval is guided by the frame elements identified in the claim understanding phase. We have manually defined the frame elements used for evidence retrieval for the frames studied in our dataset in Table 6 in the Appendix. These frame elements act as filtering conditions for querying the appropriate tables or data sources. For example, in Figure 1, the Dimension frame element is used to query the relevant table for life expectancy statistics. By focusing on specific spans of the claim (i.e., the Dimension frame element “life expectancy”), this method minimizes confusion and enhances retrieval accuracy.
3.3 Fact Verification
The final step in our pipeline involves assessing the consistency between the retrieved evidence and the original claim. Verifying the truthfulness of a claim requires not only synthesizing the retrieved evidence but also understanding how it relates to the claim and its underlying intent. For instance, in Figure 1, the claim states that Japan has the second-best life expectancy, while the data shows it actually has the highest. Although the claim is not factually precise, its core premise—that Japan has an exceptionally high life expectancy—remains valid. A robust fact verification model should ideally classify such a claim as true or mostly true. Modern LLMs are well-suited for capturing such nuances, which is why we adopt an LLM-based approach for this step.
While fine-tuned approaches have shown slightly better performance on table-based fact-checking Ye et al. (2023), we employ an LLM-based fact verification approach due to their ability to handle novel inputs without needing to retrain and for their ability to provide natural language explanations of predictions. Importantly, since fact verification is not the primary focus of our work, we keep this component model-agnostic. That is, our system can accommodate any fact verification model that accepts structured evidence and a claim as input. This design choice allows for easy substitution with future state-of-the-art models as they become available.
@@ 4.1 Voting Records
We compiled a dataset of official U.S. congressional voting records from the Congress GitHub Repository. [^3] https://github.com/unitedstates/congress The dataset includes 342,466 bills from the 93rd to the 117th Congress. It also contains biographical information for 12,677 unique members of Congress dating back to the 1st Congress. However, comprehensive roll call voting records are only available from the 101st to the 117th Congress, totaling 7,195,798 individual votes across 22,447 roll calls. Each member of Congress cast an average of 4,230 votes during this period. Accordingly, our analysis is restricted to data from the 101st Congress onward, which reflects the full span of available voting records.
@@
To verify voting records, mapping the Agent FE to the correct member of Congress is essential. We use database queries to match congress members whose names are similar to the words in the Agent FE (details in Appendix C.4). In cases of name ambiguity we default to the more recent member.
This stage presents several challenges. Claims often use nicknames, such as “Sleepy Joe” for Joe Biden or “Meatball Ron” for Ron DeSantis. To address this, we extracted two lists of political nicknames from Wikipedia Wikipedia contributors (2024a, b) to map nicknames to their corresponding congress members. Similarly, some members go by shortened or preferred names, such as “Joe Biden” instead of “Joseph Biden” or “Ted Cruz” instead of “Rafael Edward Cruz”. We handle this by utilizing a list 555 https://github.com/carltonnorthern/nicknames of congress members’ preferred names, supplemented by common alternatives.
@@
Identifying the correct bill based on the extracted Issue FE is challenging because the Issue FE can refer to various types of information, such as abstract topics (e.g., “gun control”), specific actions or bills (e.g., the “Inflation Reduction Act of 2022”), or outcomes of legislation (e.g., “preventing women from getting abortions”). Additionally, bills often do not include the colloquial terms commonly used to describe them. For example, the “STOP School Violence Act of 2018” may be informally described as expanding access to guns in schools, even though the bill itself does not use this phrasing. Because of these challenges, keyword-based search is insufficient for accurate evidence retrieval. To address this, we employ an asymmetric semantic similarity model, as described in Section 5.4, enabling us to identify bills that are semantically similar to the Issue FE, even when the style of language used in the claim and the bill differs significantly.
@@
Determining whether a claim is refuted or supported by evidence presents several challenges. First, the system cannot rely solely on the vote (Yea or Nay) and the Position FE (for or against).  [^6] Note that Position FE is extracted from claims by our frame semantic parsing model. However, only Agent and Issue FEs are used for evidence retrieval in voting-related claims. Claims may be made without a Position FE, and the relationship between the vote on a bill and the claim’s Position FE can differ. For instance, a Yea vote on a bill may not indicate support for the claimed Issue. Consider the claim, “DeSantis voted against allowing abortions,” in conjunction with a Yea vote on a bill that bans abortion. Here, the vote supports the claim despite the discrepancy between the Position FE (against) and the vote on the bill (for/yea). Second, assessing whether a claim is supported or refuted by a bill vote requires an understanding of the bill itself and its implications. To address these concerns, we instruct a large language model (LLM) with these challenges in mind using the prompt in Appendix E.2.